OpenStudy (anonymous):

express using a positive exponent h-^4

OpenStudy (anonymous):

Recall that anytime you have a negative exponent, you can make it positive by 'moving' it to the other side of the fraction bar. \[x^{-3} = \frac{1}{x^3}\]\[\frac{1}{a^{-2}} = a^2\]

OpenStudy (anonymous):

so it a^2?? wouldnt u use the h

OpenStudy (anonymous):

I wasn't giving you the answer. I was showing you how to solve it. Use what I said and apply it to your problem.

OpenStudy (anonymous):

ok thank u I was just having trouble havent took math in 15 years

OpenStudy (anonymous):

Been there, done that.

OpenStudy (anonymous):

Went back to school after 15 years and had to re-do all of algebra & trig before I could move on to calculus, etc.

OpenStudy (anonymous):

I know it is nerve wracking isn't it especially when I am a mom of 3.. OMG!!! but thank goodness this is my last math class for my associates degree.

OpenStudy (anonymous):

Did you figure out the answer?

OpenStudy (anonymous):

I think it is h-4.. IDk

OpenStudy (anonymous):

No no. You started with this: \[h^{-4}\] Now, do you see one of my examples that looks like that?

OpenStudy (anonymous):

1/h4

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

Now if you have other problems like this, they should be quite easy follow the rules I gave above.